OpenStudy (anonymous):

Summation of k=4 to 20 of k^2

OpenStudy (anonymous):

\[\sum_{4}^{20} k^2\]

OpenStudy (anonymous):

4^2 + 5^2 + .... + 20^2

OpenStudy (anonymous):

Is there a way of not expanding the entire thing?

OpenStudy (anonymous):

are you in calculus? like calc 2. are you guys doing series.

OpenStudy (anonymous):

Yes. I am in Calculus BC.

OpenStudy (anonymous):

i haven't taken it in a while so I don't remember if there is a special method to do that.

OpenStudy (anonymous):

Hmm. Ok then, thanks anyway. :)

OpenStudy (anonymous):

@Algebraic! @calculusfunctions @eseidl @radar Sorry to bother you guys, but I need to teach my class how to do this tomorrow, so any help would be greatly appreciated. :)

OpenStudy (anonymous):

There is a formula to do the sum of the squares of the first n natural numbers. It is:\[\frac{n^3}{3}+\frac{n^2}{2}+\frac{n}{6}\]you could calculate the sum of the first 25 natural numbers, taking n=25. This gives you the sum from n=0 to n=25. To get the sum from 4 to 25 you could subtract the sum you get when n=4 from the first number.

OpenStudy (anonymous):

There is a formula for the sum of the first n squares. If you write the question as:\[\sum_{k=4}^{20}k^2=\sum_{k=1}^{20}k^2-\sum_{k=1}^{3}k^2\]you could apply the formula:\[\sum_{k=1}^{n}k^2=\frac{n(n+1)(2n+1)}{6}\]

OpenStudy (anonymous):

eeer sorry, meant n=20.

OpenStudy (anonymous):

@joemath314159 , why would the top bound be 3? Wouldn't it be four?

OpenStudy (anonymous):

btw, my formula and @joemath314159 are the same thing

OpenStudy (anonymous):

Yeah, one if just simplifies more, right?

OpenStudy (anonymous):

*is ... simplified

OpenStudy (anonymous):

yeah I thought the top bound would be 4 but I could be wrong

OpenStudy (anonymous):

Because you want: 4^2+5^2+6^2+...+19^2+20^2 which is the same as: (1^2+2^2+3^2+4^2+...+19^2+20^2)-(1^2+2^2+3^2). You only want to subtract the first 3 terms so you dont accidentally get rid of the 4.

OpenStudy (anonymous):

OH, ok, that makes sense, thank you! Both of you! :D

OpenStudy (anonymous):

no, @joemath314159 must be correct. example, the sum from n=2 to n=3 would be 2^2+3^2=13 or, \[\frac{27}{3}+\frac{9}{2}+\frac{3}{6}-1=5+9-1=13\]

OpenStudy (anonymous):

Yes, he is. :) Thank you

OpenStudy (anonymous):

no problem.